OpenStudy (anonymous):

Let f(x) = { 4-x^2 if x is less tahn or = to 2 ; x-1 if x is greater than 2} a) find lim as x appraches 2 from the left f(x) and lim as x appraches 2 from teh right f(x)

OpenStudy (anonymous):

\[a) \lim_{x \rightarrow 2-} f(x) and \lim_{x \rightarrow 2+} f(x)\]

OpenStudy (anonymous):

a) As you approach it from the left, you put in values x<2, so stick with the first branch, so you get 4-2^2=0. As you approach from the right, you put in values x>2, so you get 6-2=4.

OpenStudy (anonymous):

b) does \[\lim_{x \rightarrow 2} f(x) \] exist ? c) what would te hgraph of f look like?

OpenStudy (anonymous):

b) No

OpenStudy (anonymous):

c) It makes a jump from 0 to 4 at x=2.

OpenStudy (anonymous):

how come b does not exist

OpenStudy (anonymous):

because left side and right side limits don't match up

OpenStudy (anonymous):

oh yeah i see 0 and 4 dont match

OpenStudy (ybarrap):

lim f(x) x -> 2_ = lim 4 - x^2 = lim 4 - lim x^2 = lim 4 - (lim x)^2 = 4 - 2^2 = 4 -4 = 0 lim f(x) x-> 2+ = lim (x-1) = lim x - lim(1) = 2 - 1 = 1 lim 2_ not equal to lim 2+ since 0 <> 1, therefore lim x x -> 2 does not exist.

OpenStudy (anonymous):

thanks alot... =)

OpenStudy (ybarrap):

np